# L4DR: LiDAR-4DRadar Fusion for Weather-Robust 3D Object Detection

### v5

@@ Abstract
LiDAR-based 3D object detection is crucial for autonomous driving. However, due to the quality deterioration of LiDAR point clouds, it suffers from performance degradation in adverse weather conditions. Fusing LiDAR with the weather-robust 4D radar sensor is expected to solve this problem; however, it faces challenges of significant differences in terms of data quality and the degree of degradation in adverse weather. To address these issues, we introduce L4DR, a weather-robust 3D object detection method that effectively achieves LiDAR and 4D Radar fusion. Our L4DR proposes Multi-Modal Encoding (MME) and Foreground-Aware Denoising (FAD) modules to reconcile sensor gaps, which is the first exploration of the complementarity of early fusion between LiDAR and 4D radar. Additionally, we design an Inter-Modal and Intra-Modal ({IM}2) parallel feature extraction backbone coupled with a Multi-Scale Gated Fusion (MSGF) module to counteract the varying degrees of sensor degradation under adverse weather conditions. Experimental evaluation on a VoD dataset with simulated fog proves that L4DR is more adaptable to changing weather conditions. It delivers a significant performance increase under different fog levels, improving the 3D mAP by up to 20.0% over the traditional LiDAR-only approach. Moreover, the results on the K-Radar dataset validate the consistent performance improvement of L4DR in real-world adverse weather conditions. The code can be publicly available at: https://github.com/ylwhxht/L4DR.
@@ Introduction
3D object detection is a fundamental vision task of unmanned platforms, extensively utilized in applications such as intelligent robot navigation (Ghita et al. 2024; Xu et al. 2022) and autonomous driving (Bijelic et al. 2020). For example, Full Driving Automation (FDA, Level 5) relies on weather-robust 3D object detection, which provides precise 3D bounding boxes even under various challenging adverse weather conditions (Qian et al. 2021). Owing to the high resolution and strong interference resistance of LiDAR sensors, LiDAR-based 3D object detection has emerged as a mainstream area of research (Yan et al. 2023; Wu et al. 2024; Xia et al. 2024). However, LiDAR sensors exhibit considerable sensitivity to weather conditions. In adverse scenarios, the scanned LiDAR point clouds suffer from substantial degradation and increased noise (Huang et al. 2024; Hahner et al. 2021). This degradation negatively impacts 3D detectors, compromising the reliability of autonomous perception systems.
@@
Despite progress, the significant data quality disparities and varying sensor degradation in adverse weather conditions remain largely unaddressed. As shown in Fig.2 (a), a primary challenge arises from the significant quality disparity between the LiDAR and 4D radar sensors. A second challenge pertains to varying sensor degradation under adverse weather conditions. Fig.2 (b) shows LiDAR sensors undergo severe data degradation in adverse weather. Conversely, the data quality decrease in 4D radar is significantly lower (Han et al. 2023; Sun and Zhang 2021) (Details are given in supplementary material). The varying degradation of sensors leads to fluctuations in the fused features, resulting in difficulty for the model to maintain high performance under adverse weather conditions. Therefore, developing a weather-robust 3D object detector requires overcoming the challenges of significant quality disparities and varying sensor degradation.
@@
LiDAR sensors may undergo degradation under adverse weather conditions. Physics-based simulations (Teufel et al. 2022; Hahner et al. 2022, 2021; Kilic et al. 2021) have been explored to reproduce point clouds under adverse weather to alleviate the issue of data scarcity. (Charron, Phillips, and Waslander 2018; Heinzler et al. 2020) utilized the DROR, DSOR, or convolutional neural networks (CNNs) to classify and filter LiDAR noise points. (Xu et al. 2021) designed a general completion framework that addresses the problem of domain adaptation across different weather conditions. (Huang et al. 2024) designed a general knowledge distillation framework that transfers sunny performance to rainy performance. However, these methods primarily rely on single-LiDAR modal data, which will be constrained by the decline in the quality of LiDAR under adverse weather conditions.
@@
MVDNet (Qian et al. 2021) has designed a framework for fusing LiDAR and radar. ST-MVDNet (Li et al. 2022) and ST-MVDNet++ (Li, O’Toole, and Kitani 2023) incorporate a self-training teacher-student to MVDNet to enhance the model. Bi-LRFusion (Wang et al. 2023b) framework employs a bidirectional fusion strategy to improve dynamic object detection performance. However, these studies only focus on 3D radar and LiDAR. As research progresses, the newest studies continue to drive the development of LiDAR-4D radar fusion. M2-Fusion (Wang et al. 2022a), InterFusion (Wang et al. 2022b), and 3D-LRF (Chae, Kim, and Yoon 2024) explore LiDAR and 4D radar fusion. However, these methods have not considered and overcome the challenges of fusing 4D radar and LiDAR under adverse weather conditions.
@@
As illustrated in Figure 5, we innovate the traditional unimodal Pillars coding into multimodal Pillars coding to perform initial fusion at the data level, extracting richer information for subsequent feature processing. Firstly, referring (Lang et al. 2019), we encode the LiDAR point cloud into a pillar set $P^{l}=\{p^{l}_{i}\}_{i=1}^{N}$. Each LiDAR point $p^{l}_{(i,j)}$ in pillar $p^{l}_{i}$ encoded using encoding feature $\boldsymbol{f}^{l}_{(i,j)}$ as
@@
where $\boldsymbol{\mathcal{X}}^{l}=[x^{l},y^{l},z^{l}]$ is the coordinate of the LiDAR point, $\boldsymbol{\mathcal{Y}}^{l}_{cl}$ denotes the distance from the LiDAR point to the arithmetic mean of all LiDAR points in the pillar, $\boldsymbol{\mathcal{Z}}^{l}$ denotes the (horizontal) offset from the pillar center in $x,y$ coordinates, and $\lambda$ is the reflectance. Similarly, each 4D radar point $p^{r}_{(i,j)}$ in pillar $p^{r}_{i}$ encoded using encoding feature $\boldsymbol{f}^{r}_{(i,j)}$ as
$\boldsymbol{f}^{r}_{(i,j)}=[\boldsymbol{\mathcal{X}}^{r},\boldsymbol{\mathcal{Y}}^{r}_{cr},\boldsymbol{\mathcal{Z}}^{r},\boldsymbol{\mathcal{V}},\Omega],$ (2)
where $\boldsymbol{\mathcal{X}}^{r}$, $\boldsymbol{\mathcal{Y}}^{r}_{cr}$, and $\boldsymbol{\mathcal{Z}}^{r}$ are similar in meaning to those in Eq.1. $\boldsymbol{\mathcal{V}}=[\mathcal{V}_{x},\mathcal{V}_{y}]$ is Doppler information along each axis, and $\Omega$ is Radar Cross-Section (RCS).
We then perform cross-modal feature propagation for LiDAR pillar encoding features $\boldsymbol{f}^{l}_{(i,j)}$ and radar pillar encoding features $\boldsymbol{f}^{r}_{(i,j)}$ that occupy the same coordinates. The fused LiDAR pillar encoding features $\widehat{\boldsymbol{f}}^{l}_{(i,j)}$ and 4D radar points $\widehat{\boldsymbol{f}}^{r}_{(i,j)}$ are obtained by fusing $\boldsymbol{f}^{l}_{(i,j)}$ and $\boldsymbol{f}^{r}_{(i,j)}$ as follows:
@@
$\enspace\mathrm{and}\enspace\widehat{\boldsymbol{f}}^{r}_{(i,j)}=[\boldsymbol{\mathcal{X}}^{r},\boldsymbol{\mathcal{Y}}^{r}_{cl},\boldsymbol{\mathcal{Y}}^{r}_{cr},\boldsymbol{\mathcal{Z}}^{r},\overline{\lambda},\boldsymbol{\mathcal{V}},\Omega],$
@@
Loss Function and Training Strategy.
@@
The VoD dataset (Palffy et al. 2022) contains 8693 frames of 64-line LiDAR, camera, and 4D radar data. Following the official partition, we divided the dataset into a training and validation set with 5139 and 1296 frames. All of the methods used the official radar with 5 scans accumulation and single frame LiDAR. Meanwhile, to explore the performance under different fog intensities, following a series of previous work (Qian et al. 2021; Li et al. 2022), we similarly performed fog simulations (Hahner et al. 2021) (with fog level $\mathcal{L}$ from 0 to 4, fog density ($\alpha$) = [0.00, 0.03, 0.06, 0.10, 0.20]) on the VoD dataset and kept 4D Radar unchanged to simulate various weather conditions. We named it the Vod-Fog dataset in the following. Noteworthy, we used two evaluation metrics groups on the VoD dataset. The VoD official metrics are used to compare the results reported by previous state-of-the-art methods. The KITTI official metrics are used to demonstrate and analyze the performance of ”easy”, ”moderate” and ”hard” objects with different difficulties under foggy weather.
@@
Following 3D-LRF (Chae, Kim, and Yoon 2024), we compare our L4DR with different modality-based 3D object detection methods: PointPillars (Lang et al. 2019), RTNH (Paek, KONG, and Wijaya 2022), InterFusion (Wang et al. 2022b) and 3D-LRF (Chae, Kim, and Yoon 2024). The results in Table 1 highlight the superior performance of our L4DR model on the K-Radar dataset. Our L4DR model surpasses 3D-LRF by 8.3% in total $AP_{3D}$. This demonstrates that compared to previous fusion, our method utilizes the advantages of LiDAR and 4D Radar more effectively. We observed that all methods perform better in many adverse weather conditions (e.g., Overcast, Fog, etc.) than in normal weather. A possible reason is the distribution differences of labeled objects in different weather conditions. We discuss this counter-intuitive phenomenon in detail in the supplementary material, as well as other valuable results, such as different IoU thresholds and new version labels, for a more comprehensive comparison. Note that we only compare 3D-LRF on the K-Radar dataset because its code is not open-sourced, the results of 3D-LRF are available on the K-Radar dataset only.
@@
Comparison with other feature fusion.
We compared different multi-modal feature fusion blocks, including basic concatenation (Concat.) and various attention-based methods such as Transformer-based (Vaswani et al. 2017) Cross-Modal Attention (Cross-Attn.) and Self-Attention (Self-Attn.), SE Block (Hu, Shen, and Sun 2018), and CBAM Block (Woo et al. 2018), see supplementary material for detailed fusion implements. Experimental results (Table 5) show that while attention mechanisms outperform concatenation to some extent, they do not effectively address the challenge of fluctuating features under varying weather conditions. In contrast, our proposed MSGF method, focusing on significant features of LiDAR and 4D radar, achieves superior performance and robustness under different weather.
@@
To explore the 4D radar denoising ability of FAD, we tested different $\lambda$ in the inference stage, as shown in Table 6. With hyperparameter $\lambda$ increase, recall gradually decreases, and denoise rate, mIoU, and PA gradually increase. We find that the performance of the 3D object detection is not excellent when $\lambda$ is relatively high (0.5). The reason is missing foreground points is more unacceptable than retaining some noise for object detection. Additionally, we also discuss the effect of $\lambda$ on the detection performance in the supplementary material.
@@ Conclusion
In this paper, we analyzed the challenges of fusing LiDAR and 4D radar in adverse weather and proposed L4DR, an effective LiDAR and 4D radar fusion method. We provide an innovative and feasible solution for achieving weather-robust outdoor 3D object detection in various weather conditions. Our experiments on VoD and K-Radar datasets have demonstrated the effectiveness and superiority of our method in various simulated fog levels and real-world adverse weather. In summary, our proposed L4DR fusion method not only offers a promising solution for robust outdoor 3D object detection in adverse weather conditions but also sets a new benchmark for performance and robustness compared to existing fusion techniques, paving the way for enhanced safety and reliability in autonomous driving and other applications.

### v6

@@ Abstract
LiDAR-based 3D object detection is crucial for autonomous driving. However, due to the quality deterioration of LiDAR point clouds, it suffers from performance degradation in adverse weather conditions. Fusing LiDAR with the weather-robust 4D radar sensor is expected to solve this problem; however, it faces challenges of significant differences in terms of data quality and the degree of degradation in adverse weather. To address these issues, we introduce L4DR, a weather-robust 3D object detection method that effectively achieves LiDAR and 4D Radar fusion. Our L4DR proposes Multi-Modal Encoding (MME) and Foreground-Aware Denoising (FAD) modules to reconcile sensor gaps, which is the first exploration of the complementarity of early fusion between LiDAR and 4D radar. Additionally, we design an Inter-Modal and Intra-Modal ({IM}2) parallel feature extraction backbone coupled with a Multi-Scale Gated Fusion (MSGF) module to counteract the varying degrees of sensor degradation under adverse weather conditions. Experimental evaluation on a VoD dataset with simulated fog proves that L4DR is more adaptable to changing weather conditions. It delivers a significant performance increase under different fog levels, improving the 3D mAP by up to 20.0% over the traditional LiDAR-only approach. Moreover, the results on the K-Radar dataset validate the consistent performance improvement of L4DR in real-world adverse weather conditions.
Codehttps://github.com/ylwhxht/L4DR
Extended versionhttps://arxiv.org/abs/2408.03677
@@ Introduction
3D object detection is a fundamental vision task of unmanned platforms, extensively utilized in applications such as intelligent robot navigation (Ghita et al. 2024; Xu et al. 2022) and autonomous driving (Bijelic et al. 2020). For example, Full Driving Automation (FDA, Level 5) relies on weather-robust 3D object detection, which provides precise 3D bounding boxes even under various challenging adverse weather conditions (Qian et al. 2021). Owing to the high resolution and strong interference resistance of LiDAR sensors, LiDAR-based 3D object detection has emerged as a mainstream area of research (Wu et al. 2024; Xia et al. 2024; Deng et al. 2025). However, LiDAR sensors exhibit considerable sensitivity to weather conditions. In adverse scenarios, the scanned LiDAR point clouds suffer from substantial degradation and increased noise (Huang et al. 2024a; Hahner et al. 2021). This degradation negatively impacts 3D detectors, compromising the reliability of autonomous perception systems.
@@
Despite progress, the significant data quality disparities and varying sensor degradation in adverse weather conditions remain largely unaddressed. As shown in Fig.2 (a), a primary challenge arises from the significant quality disparity between the LiDAR and 4D radar sensors. A second challenge pertains to varying sensor degradation under adverse weather conditions. Fig.2 (b) shows LiDAR sensors undergo severe data degradation in adverse weather. Conversely, the data quality decrease in 4D radar is significantly lower (Han et al. 2023; Sun and Zhang 2021) (Details are given in extended version (Huang et al. 2024b)). The varying degradation of sensors leads to fluctuations in the fused features, resulting in difficulty for the model to maintain high performance under adverse weather conditions. Therefore, developing a weather-robust 3D object detector requires overcoming the challenges of significant quality disparities and varying degradation.
@@
LiDAR sensors may undergo degradation under adverse weather conditions. Physics-based weather simulations (Hahner et al. 2022, 2021; Kilic et al. 2021) have been explored to produce point clouds under adverse weather to alleviate data scarcity. (Charron, Phillips, and Waslander 2018; Heinzler et al. 2020) utilized the DROR, DSOR, or CNN to classify and filter LiDAR noise points. (Xu et al. 2021) designed a general completion framework that addresses the domain gaps across different weather conditions. (Huang et al. 2024a) designed a general knowledge distillation framework that transfers sunny performance to rainy performance. However, these methods primarily rely on single-LiDAR modal data, which will be constrained by the decline in the quality of LiDAR under adverse weather conditions.
@@
MVDNet (Qian et al. 2021) has designed a framework for fusing LiDAR and radar. ST-MVDNet (Li et al. 2022) and ST-MVDNet++ (Li, O’Toole, and Kitani 2023) incorporate a self-training teacher-student to MVDNet to enhance the model. Bi-LRFusion (Wang et al. 2023b) employs a bidirectional fusion strategy to improve dynamic object detection performance. However, these studies only focus on 3D radar and LiDAR. As research progresses, the newest studies continue to drive the development of LiDAR-4D radar fusion. M2-Fusion (Wang et al. 2022a), InterFusion (Wang et al. 2022b), and 3D-LRF (Chae, Kim, and Yoon 2024) explore LiDAR and 4D radar fusion. However, these methods have not considered and overcome the challenges of fusing 4D radar and LiDAR under adverse weather conditions.
@@
As illustrated in Figure 5, we innovate the traditional unimodal Pillars coding into multimodal Pillars coding to perform initial fusion at the data level, extracting richer information for subsequent feature processing. Firstly, referring (Lang et al. 2019), we encode the LiDAR point cloud into a pillar set $\mathcal{H}^{l}=\{h^{l}_{i}\}_{i=1}^{N}$. Each LiDAR point $p^{l}_{(i,j)}$ in pillar $h^{l}_{i}$ encoded using encoding feature $\boldsymbol{f}^{l}_{(i,j)}$ as
@@
where $\boldsymbol{\mathcal{X}}^{l}=[x^{l},y^{l},z^{l}]$ is the coordinate of the LiDAR point, $\boldsymbol{\mathcal{Y}}^{l}_{cl}$ denotes the distance from the LiDAR point to the arithmetic mean of all LiDAR points in the pillar, $\boldsymbol{\mathcal{Z}}^{l}$ denotes the (horizontal) offset from the pillar center in $x,y$ coordinates, and $\lambda$ is the reflectance. Similarly, each 4D radar point $p^{r}_{(i,k)}$ in pillar $h^{r}_{i}$ encoded using encoding feature $\boldsymbol{f}^{r}_{(i,k)}$ as
$\boldsymbol{f}^{r}_{(i,k)}=[\boldsymbol{\mathcal{X}}^{r},\boldsymbol{\mathcal{Y}}^{r}_{cr},\boldsymbol{\mathcal{Z}}^{r},\boldsymbol{\mathcal{V}},\Omega],$ (2)
where $\boldsymbol{\mathcal{X}}^{r}$, $\boldsymbol{\mathcal{Y}}^{r}_{cr}$, and $\boldsymbol{\mathcal{Z}}^{r}$ are similar in meaning to those in Eq.1. $\boldsymbol{\mathcal{V}}=[\mathcal{V}_{r},\mathcal{V}_{a}]$ is the relative and absolute radial velocity of Doppler information, and $\Omega$ is Radar Cross-Section (RCS).
We then perform cross-modal feature propagation for LiDAR and radar pillar encoding features that occupy the same pillar coordinates. The fused LiDAR and 4D radar pillar encoding features $\widehat{\boldsymbol{f}}^{l}_{(i,j)}$ and $\widehat{\boldsymbol{f}}^{r}_{(i,k)}$ are obtained by fusing $\boldsymbol{f}^{l}_{(i,j)}$ and $\boldsymbol{f}^{r}_{(i,k)}$ as follows:
@@
$\enspace\mathrm{and}\enspace\widehat{\boldsymbol{f}}^{r}_{(i,k)}=[\boldsymbol{\mathcal{X}}^{r},\boldsymbol{\mathcal{Y}}^{r}_{cl},\boldsymbol{\mathcal{Y}}^{r}_{cr},\boldsymbol{\mathcal{Z}}^{r},\overline{\lambda},\boldsymbol{\mathcal{V}},\Omega],$
@@
Loss Function and Training Strategy
@@
The VoD dataset (Palffy et al. 2022) contains 8693 frames of 64-line LiDAR, camera, and 4D radar data. Following the official partition, we divided the dataset into a training and validation set with 5139 and 1296 frames. All of the methods used the official radar with 5 scans accumulation and single frame LiDAR. Meanwhile, to explore the performance under different fog intensities, following a series of previous work (Qian et al. 2021; Li et al. 2022), we similarly performed fog simulations (Hahner et al. 2021) (with fog level $\mathcal{L}$ from 0 to 4, fog density ($\alpha$) = [0.00, 0.03, 0.06, 0.10, 0.20]) on the VoD dataset and kept 4D Radar unchanged to simulate various weather conditions. We named it the Vod-Fog dataset in the following. Noteworthy, we used two evaluation metrics groups on the VoD dataset. The VoD official metrics are used to compare with previous state-of-the-art methods. The KITTI official metrics are used to analyze the performance of objects with different difficulties.
@@
Following 3D-LRF (Chae, Kim, and Yoon 2024), we compare our L4DR with different modality-based 3D object detection methods: PointPillars (Lang et al. 2019), RTNH (Paek, KONG, and Wijaya 2022), InterFusion (Wang et al. 2022b) and 3D-LRF (Chae, Kim, and Yoon 2024). The results in Table 1 highlight the superior performance of our L4DR model on the K-Radar dataset. Our L4DR model surpasses 3D-LRF by 8.3% in total $AP_{3D}$. This demonstrates that compared to previous fusion, our method utilizes the advantages of LiDAR and 4D Radar more effectively. We observed that all methods perform better in many adverse weather conditions (e.g., Overcast, Fog, etc.) than in normal weather. A possible reason is the distribution differences of labeled objects in different weather conditions. We discuss this counter-intuitive phenomenon in the extended version and other valuable results, such as different IoU thresholds and version labels.
@@
Comparison with other feature fusion.
We compared different multi-modal feature fusion blocks, including basic concatenation (Concat.) and various attention-based methods such as Transformer-based (Vaswani et al. 2017) Cross-Modal Attention (Cross-Attn.) and Self-Attention (Self-Attn.), SE Block (Hu, Shen, and Sun 2018), and CBAM Block (Woo et al. 2018), see extended version for detailed fusion implements. Experimental results (Table 5) show that while attention mechanisms outperform concatenation to some extent, they do not effectively address the challenge of fluctuating features under varying weather conditions. In contrast, our proposed MSGF method, focusing on significant features of LiDAR and 4D radar, achieves superior performance and robustness under different weather.
@@
To explore the 4D radar denoising ability of FAD, we tested different $\lambda$ in the inference stage, as shown in Table 6. With hyperparameter $\lambda$ increase, recall gradually decreases, and denoise rate, mIoU, and PA gradually increase. We find that the performance of the 3D object detection is not excellent when $\lambda$ is relatively high (0.5). The reason is missing foreground points is more unacceptable than retaining some noise for object detection. Additionally, we also discuss the effect of $\lambda$ on the detection performance in the extended version.
@@ Conclusion
In this paper, we analyzed the challenges of fusing LiDAR and 4D radar in adverse weather and proposed L4DR, an effective LiDAR and 4D radar fusion method. We provide an innovative and feasible solution for achieving weather-robust outdoor 3D object detection in various weather conditions. Our experiments on VoD and K-Radar datasets have demonstrated the effectiveness and superiority of our method in various simulated fog levels and real-world adverse weather. In summary, our proposed L4DR paves the way for enhanced reliability in autonomous driving and other applications.
@@
Acknowledgements
This work was supported in part by the National Natural Science Foundation of China (No.62171393), and the Fundamental Research Funds for the Central Universities (No.20720220064).